M^2-8m-37=-4
I'm not even sure how to start I think I am supposed to complete the square
^2 means the number is squared

Answer by mananth(16946)

m^2-8m-37=-4
m^2-8m= 37-4
m^2-8m=33
m^2-8m+16=33+16
(m-4)^2=49
take the square root
m-4= +/-7
m=11 or -3